OpenStudy (punderson):

Choose the equation of the horizontal line that passes through the point (1, -5). @decentnabeel y = -5 x = 1 y = 1 x = -5

OpenStudy (punderson):

Choose the equation of the line passing through the point (-4, -2) and parallel to y = x + 2. (1 point) y = 4x - 1 y = 4x - 3 y = x - 3 y = x - 1

OpenStudy (punderson):

Choose the equation of the line passing through the point (-4, -2) and perpendicular to y = -x + 6. (1 point) y = -x + 6 y = x - 6 y = -x - 2 y = x + 2

OpenStudy (punderson):

The equation of a line is shown below: y = - x - 24 What is the equation of a line which is perpendicular to this line and passes through (1, 27) (1 point) y = 3x - 24 y = 7x - 24 y = 7x + 24 y = 3x + 24

alones (alones):

Umm can you post each question one at a time?